OpenStudy (mathguy5):

You invested $22,000 in two accounts paying 6% and 8% annual interest , respectively , if the total interest earned for the year is was $1700, how much was invested for each rate. At 6%= At 8%=

OpenStudy (kinged):

Start with 0.06x + 0.08y = 1700 (Given that x is for first account and y is for second account.)

OpenStudy (mathguy5):

i got 0.6+1760=0.8=1700

OpenStudy (kinged):

How does 0.8=1700?

OpenStudy (mathguy5):

i meant minus sorry

OpenStudy (kinged):

Just discard that equation from your mind. Think this way: The $22,000 were separated into two accounts (with amounts x and y). These amounts yield 6% for x and 8% for y (That's 0.06 and 0.08, respectively).

OpenStudy (mathguy5):

0.6x0.8(22,000-x)=1700

OpenStudy (kinged):



OpenStudy (kinged):

Discard the $22,000 from the equation for now.

OpenStudy (mathguy5):

I got 1000 and 700 as the 2 final answers

OpenStudy (kinged):

But these numbers don't add up to $22,000 (the number you invested in both).

OpenStudy (mathguy5):

I think its 2000 and 200 since x gives 2000 and the remaining has to be 200.

OpenStudy (mathguy5):

i mean 20,000 and 200

OpenStudy (kinged):

20,000+200=20,200 not 22,000

OpenStudy (mathguy5):

thanks. I mean 20,000 and 2000

OpenStudy (mathstudent55):

Why are you guessing? Write a system of equations, solve it, and you'll know for sure. So far you haven't guessed the correct answer yet.

OpenStudy (mathguy5):

after reviewing i got y=0.75x+21250

OpenStudy (kinged):

Use these two equations: The total sum of investments in both accounts which is x+y=? then make an equation for the yield which is 00.6x+0.08y=?

OpenStudy (kinged):

Complete these steps so that we can go on.